# Tailoring transport in quantum spin chains via disorder and collisions

### v3

@@ III.2 Multiple excitations
To shape our understanding of the system’s dynamics, we analyze the case of multiple excitations. This introduces additional complexity since interactions between the excitations mediated by anisotropy, disorder, and noise, play a critical role in determining the transport and localization behavior. We systematically analyze how these interactions evolve over time and their dependence on the collision rate $r_{c}$, the disorder range constant $h$, the shape parameter $\nu$, and the anisotropy $\Delta$. Before proceeding with this program, following the outcomes of [16], we first analyze how the anisotropy affects the delocalization behavior. This will allow us to identify the most interesting regimes to zoom in. For more than one excitation IPR is no longer suited and, following [16], we resort to the Inverse Ergodicity Ratio (IER).
III.2.1 Effect of anisotropy for two excitations
To capture these effects, we consider two distinct initial configurations for just two excitations: adjacent to each other and separated by two spins, respectively.
@@ III.2.1 Effect of anisotropy for two excitations
Figure 3: Effect of anisotropy: the case of two excitations. Inverse Ergodicity ratio (IER) with associated magnetization, plateau length and delocalization time $\tau$ for a spin chain of $N=20$ sites and final time of simulations $tJ=30$. (a) IER vs time $tJ$ for different collision rates $r_{c}$ with anisotropy strength $\Delta=2.5$. Inset: magnetization map over sites $i$ and time $tJ$, showing the spreading of the two excitations over time. The two neighbouring excitations are initially separated by two spins with very homogeneous collisions over time ($\nu=100$) and high disorder range with $h=10$. The IER has the same trend of the IPR and we focus here on the collision rate regime $0\leq r_{c}\leq 1$, proved to be more interesting due to the formation of plateaus. For lower collision rates ($r_{c}\ll 1$) the systems remains localized (IER closer to 1), while increasing $r_{c}$ leads to a faster delocalization (IER decaying to its limit). (b) IER vs $tJr_{c}$ varying the collision rate parameter $0\leq r_{c}\leq 1$, at fixed high time-homogeneity ($\nu=100$), high disorder strength ($h=10$), and in the presence of anisotropy, considering two excitations initially separated by two spins. Curves with progressively increasing thickness refer to values of $r_{c}$ increasing from 0.20 to 1.00 by 0.05 each. We see that all curves collapse in the region where plateaus are present since they are naturally shifted with respect to each other. (c) 3D plot of plateau length $L$ in terms of tJ vs disorder range $h\in[0,10]$ and collision rate $r_{c}\in[0,1]$. (d) 3D plot of the complete delocalization time $\tau$ vs $h\in[0,10]$ and $r_{c}\in[0,10]$. The plane $r_{c}=h$ separates high $\tau$ regions from the small $\tau$ ones. In (c) and (d) different colors and shapes for the symbols represent different levels of collision time-homogeneity (colors) and anisotropy (shapes) as in the legend. Also, we show data from two excitations initially injected both next to each other and separated by two spins. The simulations were performed with $M=250$ trajectories and $dt=0.02$.
@@
We show in Fig. 3(a) the time evolution of the IER for two excitations initially injected at a distance of two spins apart in a system of $N=20$ sites, subjected to collision rates $r_{c}$ (from 00 to $1$) at fixed temporal homogeneity $\nu=100$, anisotropy $\Delta=2.5$, and a high disorder range $h=10$. The general trend indicates that lower collision rates ($r_{c}\ll 1$) lead to prolonged localization with the IER remaining closer to 1: the system’s state is dominated by a limited number of basis configurations. As $r_{c}$ increases, the IER decreases more rapidly over time, signaling enhanced delocalization due to frequent noise-induced perturbations. For intermediate collision rates $r_{c}\sim 0.5$, we observe a gradual reduction of the IER, while for higher rates $r_{c}\geq 1$, the system rapidly approaches ergodic behavior. For completeness, we display in the inset of Fig. 3(a) the magnetization spread at fixed collision rate $r_{c}=0.1$, the other parameters being unchanged.
@@
Consistently with the single-excitation scenario, the emergence of plateaus in the IER remains evident in the presence of noise. However, we now see that the plateaus are stabilized by the anisotropy interaction term even for lower disorder (small $h$ values) when the excitations are injected next to each other (see also the difference between Fig. 7(a) and Fig. 7(b) in Appendix B). This suggests that when more excitations are present, anisotropy works to favor the system localization.
@@
In contrast, an increase in the collision rate promotes transport by breaking the pinning of the excitations, an effect facilitated by noise. This can be seen as a manifestation of stochastic resonance, in fact happening also without disorder [16]. In addition, in the case of two separated excitations at the beginning, the presence of anisotropy doesn’t modify the IER behaviour for small values of $h$.
@@
Consequently, we come to the conclusion that when the two excitations start next to each other and the presence of interaction is significant, the IER behaviour is affected even for weaker disorder. Instead, when they start separated, the impact of their interaction can be considered a sub-leading effect, not significantly affecting the IER. We notice from Fig. 7(a) that for larger collision rates, before the plateaus are well developed, the IER curves at different $r_{c}$ values tend to collapse, indicating indeed that delocalization gets progressively insensitive to collision events once these become more frequent. It is therefore natural to ask whether a similar universal-type behavior can be identified also in the regime of rare collision events. To this aim, we show the same data in Fig. 7(b) but now with time rescaled with $r_{c}$. We indeed see that all curves collapse in the region where plateaus are present. This occurs because the plateau length scales as $1/r_{c}$ and the different curves are thus naturally shifted with respect to each other.
@@
To have a complete understanding of the plateaus formation, we systematically explored how all the governing system parameters (disorder, time and space homogeneity, and anisotropy) affect the degree of localization. Our findings are summarized in Fig. 3(c)-(d), highlighting an intricate interplay in shaping the localization dynamics. Fig. 3(c) depicts in a 3D plot the plateau width as a function of the collision rate $r_{c}$ and the disorder range constant $h$, while varying the shape parameter $\nu$ and the anisotropy $\Delta$, as depicted in the legend.
@@
Overall, our results reveal that delocalization is facilitated by conditions that minimize the stabilizing effects of disorder and anisotropy. Low disorder levels ($h\ll 10$) and weak anisotropy allow noise-induced scattering to overcome any tendency toward localization, especially when combined with a higher collision rate ($1\leq r_{c}\leq 10$), which disrupts coherent trapping mechanisms. Furthermore, lower values of the shape parameter $\nu$, corresponding to temporally heterogeneous noise, contribute to more dynamic and irregular scattering events, also promoting delocalization. In contrast, higher levels of disorder ($h\geq 10$) and anisotropy, coupled with lower collision rates ($r_{c}\ll 1$), result in a pronounced localization. In this regime, disorder-induced traps dominate the dynamics, leading to extended plateaus where the system remains effectively localized over long timescales. We observe that the length of these plateaus scales inversely with the collision rate ($\sim 1/r_{c}$), consistently with the one excitation case: indeed low collision frequencies allow the system to persist in localized configurations for extended periods before noise-induced delocalization sets in.
[…]
@@ III.2.2 The full picture
We are now ready to get the full picture, extending our investigation to explore the more general case of multiple excitations within the disordered and noisy spin chain.
@@
To this aim, we shrink our spin chain to $N=8$ and inject $N_{exc}=4$ into the system, starting all next to each other to take into account the case where anisotropy plays a role. As we learned from Sec. III.2.1 indeed, we restrict the parameter range to the most significant behavior. To sum up, this is for space-homogeneous (e.g. $\nu=100$) collisions happening at low rates ($0.1\leq r_{c}\leq 5)$, high disorder range bandwidth (e.g. $h=10$) and anisotropy as large as $\Delta=2.5$.
@@
We display in Fig. 4(a) the time evolution of the Inverse Ergodicity Ratio (IER) up to enough long time $tJ=1000$, to let the system evolve and potentially reach a delocalized state. We see that for low collision rates ($r_{c}\ll 1$) combined with time-homogeneous collisions, ($\nu\gg 1$), the system is localized over a significant portion of the evolution, especially in the presence of anisotropy and high disorder ($h\geq 10$). In such conditions, the interplay of disorder, anisotropy, and noise stabilizes localized states delaying the onset of delocalization as it is zoomed in the inset. In this case, the height of the plateaus is shorter compared to the case of one and two excitations and complete delocalization is more difficult to reach. Thus, only at very late times does the IER approach its asymptotic value, indicating that the system has completely delocalized.
@@
To gain insight on the underlying spatial behavior, we look in Fig. 4(b) at the spread of magnetization over time, displayed as a color plot. We see that for low collision rates and high time-homogeneous collisions, the magnetization piles up for extended periods in specific regions of the spin chain, highlighting the system’s localized nature. Anisotropy and disorder favor the pinning of excitations to certain sites. As time progresses, however, collisions eventually disrupt these pinned configurations, leading to gradual spread of magnetization across the chain. Delocalization is however significantly delayed under these conditions, requiring long timescales for the system to be fully delocalized.
@@
Figure 4: Case of multiple excitations. Inverse Ergodicity Ratio (IER) (a) and color plot of the magnetization (b) for a spin chain of $N=8$ sites and $N_{exc}=4$ excitations starting in the middle sites, time-homogeneous collisions $\nu=100$, with high disorder range width $h=10$ and anisotropy strength $\Delta=2.5$. (a) IER vs time $tJ$ for different collision rates $r_{c}$ in the interesting regime as in the legend. (b) Spread of magnetization over time $tJ$ and sites $i$ in the interesting regime of low collision rates, here e.g. with $r_{c}=0.1$. We can overall see that low collision rates ($r_{c}\ll 1$) combined with time-homogeneous collisions ($\nu=100$) lead the system to remain localized for a significant time, prolonged also by the presence of anisotropy and high disorder. Only at very long times the IER approaches its asymptotic value, signaling full delocalization. The simulations were performed with $M=250$ trajectories and $dt=0.02$. Final time is $tJ=1000$.

### v5

@@ III.2 Multiple excitations
To shape our understanding of the system’s dynamics, we now analyze the case of multiple excitations. Their interactions mediated by anisotropy, disorder and noise introduce additional complexity. We therefore switch to the Inverse Ergodicity Ratio (IER) described in Sec. II.1 [AlessandroetAl] and systematically analyze the interplay of the various parameters ($\nu$, $r_{c}$, $h$, $\Delta$) to map out plateau formation and delocalization scaling.
III.2.1 Two excitations case
Let us step first on the case of two excitations injected in a chain subjected to noise with homogeneous collisions.
@@ III.2.1 Two excitations case
Figure 3: The case of two excitations. (a) Inverse Ergodicity Ratio (IER) vs time $tJ$ for different collision rates $r_{c}$ with anisotropy strength $\Delta=2.5$. Inset: magnetization map over sites $i$ and time $tJ$, showing the spreading of the two excitations over time. The two neighboring excitations are initially separated by two spins with very homogeneous collisions over time ($\nu=100$) and high disorder range with $h=10$. The IER has the same trend of the IPR and we focus here on the collision rate regime $0\leq r_{c}\leq 10$, proved to be more interesting due to the formation of plateaus. For lower collision rates ($r_{c}\ll 1$) the system remains localized (IER closer to 1), while increasing $r_{c}$ leads to a faster delocalization (IER decaying to its limit). (b) IER vs $tJr_{c}$ varying the collision rate parameter $0\leq r_{c}\leq 1$, at fixed high time-homogeneity ($\nu=100$), high disorder strength ($h=10$), and in the presence of anisotropy, considering two excitations initially separated by two spins. Curves with progressively increasing thickness refer to values of $r_{c}$ increasing from 0.20 to 1.00 by 0.05 each. We see that all curves collapse in the region where plateaus are present since they are naturally shifted with respect to each other. The simulations were performed for a spin chain of $N=20$ sites with two excitations initially separated by two spins with $M=250$ trajectories and timestep $dt=0.02$ and final time of simulations $tJ=30$.
Figure 4: Characterization of the plateaus. (a) 3D plot of plateau duration $D$ in terms of tJ vs disorder range $h\in[0,10]$ and collision rate $r_{c}\in[0,1]$ (see Appendix B, Fig. 10, for a 2D version). Different symbols colors and shapes distinguish different levels of collision time-homogeneity (colors) and anisotropy (shapes), as in the legend. (b) Representing all the relevant datasets with only two quantities that collect all the governing physical parameters. Semi-log plot of the dimensionless plateaus area $Z_{J}D(tJ)r_{c}$ vs the disorder power per unit collision time $P_{h}\equiv hr_{c}$, scaled with the shape parameter $\nu$ (see text for a discussion). The datasets can be identified from the legend: different symbols label the collision rates $r_{c}=0.1,0.2,0.5,0.7,0.9,1$, different colors the disorder strength values $h=0.5,1,5,10$. Notice that all the datasets collapse in two curves: in essence, no plateaus for low disorder for any other parameter (bluish symbols) and rapidly developing plateaus to a saturating value with increasing disorder power, for slow collision rates and any other parameter (yellow-reddish symbols). As already commented, the anisotropy $\Delta$ does not qualitative impact the results and it is therefore fixed at $\Delta=2.5$ (that is, other $\Delta$ values would be not distinguishable). Also, adding the data with $r_{c}>1$ would result in having an horizontal lineup of data points collapsed at vanishing plateaus area irrespective of all the other parameters, and are thus not reported. The simulations were performed for a spin chain of $N=20$ sites with two excitations initially separated by two spins, with $M=250$ trajectories, timestep $dt=0.02$, and final time of simulations $tJ=30$.
Figure 5: Study of delocalization time $\tau$ (see text, Sec. III.2). 3D plot of the complete delocalization time $\tau$ vs $h\in[0,10]$ and $r_{c}\in[0,10]$. The plane $r_{c}=h$ separates small from intermediate-to-high $\tau$ regions. Region (1) with small $\tau$ does not exhibit plateaus, while the region with plateaus occurs in two regimes: with intermediate (2) and long (3) $\tau$, corresponding to a more localized/delocalized behavior (see text, Sec. III.2). Symbols with different colors represent different levels of collision time-homogeneity as in the legend. The simulations were performed for a spin chain of $N=20$ sites with two excitations initially separated by two spins with $M=250$ trajectories and timestep $dt=0.02$ and final time of simulations $tJ=30$. The data with $\tau>30$ are those for which the system does not succeed to reach complete delocalization in that time.
@@
We show in Fig. 3(a) the time evolution of the IER for two excitations initially injected at a distance of two spins apart in a system of $N=20$ sites, subjected to collision rates $r_{c}$ (from 0 to $10$) at fixed temporal homogeneity $\nu=100$, anisotropy $\Delta=2.5$, and a high disorder range $h=10$. The general trend indicates that lower collision rates ($r_{c}\ll 1$) lead to prolonged localization with the IER remaining closer to 1: the system’s state is dominated by a limited number of basis configurations. As $r_{c}$ increases, the IER decreases more rapidly over time, signaling enhanced delocalization due to frequent noise-induced perturbations. For $r_{c}\sim 0.5$, we observe a gradual reduction of the IER, while for higher rates $r_{c}\geq 1$, the system rapidly approaches ergodic behavior. For completeness, we display in the inset of Fig. 3(a), the magnetization spread at fixed collision rate $r_{c}=0.1$, the other parameters being unchanged.
@@
Consistently with the single-excitation scenario, the emergence of plateaus in the IER remains evident in the presence of homogeneous noise with a low collision rate. The presence of anisotropy does not modify the IER trend when the two excitations start separated by two spins, thus in this case can be considered a sub-leading effect. However, the plateaus are stabilized by the anisotropy interaction term even for lower disorder (small $h$ values) when the excitations are injected next to each other (see the difference between Fig. 9(a) and Fig. 9(b) in Appendix B). This suggests that when more excitations are present, anisotropy works to favor the system localization.
@@
In contrast with the case of one excitation, an increase in the collision rate ($r_{c}=1,5,10$) promotes transport by breaking the pinning of the excitations, an effect facilitated by noise. In fact, the presence of more than one excitation leads to a rapid decrease of the IER for higher values of the collision rate. This can be seen as a manifestation of stochastic resonance, in fact happening also without disorder [AlessandroetAl].
@@
We notice from Fig. 3(a) that, before the plateaus are well developed, the IER curves at different $r_{c}$ values tend to collapse, indicating indeed that delocalization gets progressively insensitive to collision events once these become more frequent. It is therefore natural to ask whether a universal-type behavior can be identified in the regime of rare collision events. To this aim, we show the IER vs. time rescaled with $r_{c}$ for two excitations initially separated by two spins in the same spin chain with $N=20$ sites, in the case of anisotropy $\Delta=2.5$, high shape parameter $\nu=100$, high disorder range $h=10$, and varying the collision rate $r_{c}$ between 0.2 and 1 (Fig. 3(b)). We indeed see that all curves collapse in the region where plateaus are present. This occurs because the plateau duration scales as $1/r_{c}$ and the different curves are thus naturally shifted with respect to each other.
@@
To have a complete understanding of the plateaus formation, we have systematically explored how all the governing system parameters (disorder, time and space homogeneity, and anisotropy) affect the degree of localization.
To this end, we introduce the duration of the plateaus $D(tJ)$ and the jump size $Z_{J}$. $D(tJ)$ is the duration of the first plateau for each curve along the x-axis in units of $tJ$. $Z_{J}$ is the height of the first plateau for each curve.
In Fig. 4(a), we show the 3D plot of $D(tJ)$ as a function of the collision rate $r_{c}$ and the disorder range constant $h$, with the shape parameter $\nu$ and the anisotropy $\Delta$ varied as in the legend. We can see here that the data with different values of anisotropy collapse, thus from now on in this section we will present only data with $\Delta=2.5$. (For a differently expressive 2D version of the plot, we refer to Appendix B, Fig. 10 (a)). To make order in all the data, we identify physically meaningful quantities that can collect all the governing system parameters. To start with, we consider the plateaus area, that is their height, i.e. the jump size $Z_{J}$, times their dimensionless duration $D(tJ)r_{c}$: this quantity can be viewed as a sort of order parameter for the appearance of the plateaus. We then collect the remaining physical quantities in $hr_{c}\nu$, the rate of change of disorder bandwidth $h$ per unit collision time $r_{c}$, scaled with the shape parameter $\nu$: in fact, $P_{h}\equiv h\cdot r_{c}$ can be viewed as form of disorder power in collision-time units. We present in Fig. 4(b) a semi-log plot of $Z_{J}D(tJ)r_{c}$ vs. $P_{h}\nu\equiv hr_{c}\nu$. The remaining parameters are as usual: $N=20$ sites for the spin chain, two excitations initially separated by two spins, anisotropy $\Delta=2.5$, collision rates $r_{c}$ in the range between 0.1 and 1 (labeled by different shapes as in the legend), different disorder-range values $h=0.5,1,5,10$ (shown with different colors as in the legend), and different values for the shape parameter $\nu$ in the homogeneous range between 50 and 100. We see that all the data collapse in essentially two curves, distinguishable at a glance by bluish and yellow-reddish symbols. In fact, we can easily identify the region corresponding to no plateaus, i.e. with vanishing plateaus area: this corresponds to low values of disorder strength (0.5 and 1, blue and light-blue dots) and to high collision rates $r_{c}>1$ interdependently of all the other parameters. For larger values of $h>1$, instead, the plateaus rapidly develop with decreasing the collision rate $r_{c}$, up to a size that saturates to a value of $\approx 0.2$ independently of the other parameters. While this value could change with the number of excitations involved, it would be interesting to explore the degree to which it might have a universal meaning.
@@
Before proceeding further, let us summarize our findings so far. Delocalization is facilitated by conditions that minimize the stabilizing effects of disorder and anisotropy. Low disorder levels ($h\ll 10$) and weak anisotropy allow noise-induced scattering to overcome any tendency toward localization. This occurs especially when combined with a higher collision rate ($1\leq r_{c}\leq 10$), which disrupts coherent trapping mechanisms. Furthermore, lower values of the shape parameter $\nu\sim 50$ contribute to more dynamic and irregular scattering events, also promoting delocalization. In contrast, higher levels of disorder ($h\geq 10$) and anisotropy, coupled with lower collision rates ($r_{c}\ll 1$), result in a pronounced localization. In this regime, disorder-induced traps dominate the dynamics, leading to extended plateaus where the system remains effectively localized over long timescales. We observe that the duration of these plateaus scales inversely with the collision rate ($\sim 1/r_{c}$), consistently with the one excitation case: indeed, low collision frequencies allow the system to persist in localized configurations for extended periods of time, before noise-induced delocalization sets in.
To gain further insight on the system delocalization behavior, we now introduce the complete delocalization time $\tau$. This is defined as the first time (in terms of tJ) at which the IER falls below $10/dimH$ (10 times its lower bound), and after which no additional plateaus are clearly distinguishable. Our findings summarized in Fig. 5 highlight an intricate interplay in shaping the localization dynamics. In Fig. 5 we show $\tau$ in a 3D plot again versus disorder range $h$ and collision rate $r_{c}$, and for different values of time-homogeneous shape parameters $\nu$ and anisotropy $\Delta=2.5$. (For a differently expressive 2D version of the plot, we refer to Appendix B, Fig. 10 (b))
We again identify two regions: small and intermediate-to-long $\tau$. The first region (1) refers to a complete delocalization time $\tau\leq 10$, here happening for two datasets: data with low collision rates $0.1\leq r_{c}\leq 1$ combined with small level of disorder $h\leq 1$ and data with higher collision rates $r_{c}>1$. In all cases, $\tau$ is small independently of the shape parameter $\nu$ and no plateaus are sizable. In the second region pleateaus build up but in two different regimes, (2) and (3). In (2), complete delocalization occurs over a time $10\leq\tau\leq 30$, happening for higher level of disorder $h=5,10$ combined with an intermediate collision rate $1\leq r_{c}\leq 5$. Finally, in (3) complete delocalization occurs over time $\tau>30$, occurring for high disorder with $h=10$, high collision time homogeneity $\nu=100$, and small collision rates $r_{c}\leq 1$. Notice that this region, not determined in our observation window for the IER has not reached its lower bound, corresponds to the formation of the longest plateaus. Using the plane $r_{c}=h$ as a reference, we observe that for $r_{c}<h$, the complete delocalization time remains high, while for $r_{c}>h$, $\tau$ is significantly reduced. This behavior resembles the case without disorder ([AlessandroetAl]), where the complete delocalization time was indeed limited to $\tau\leq 5$.
Summing up, we can conclude how delocalization turns out to be facilitated by low levels of disorder, anisotropy, and time homogeneity, along with a higher collision rate ($1<r_{c}<10$). Instead, higher levels of disorder and anisotropy, coupled with a lower collision rate, lead to more localized behavior, with the size of the plateaus scaling as $1/r_{c}$.
@@ III.2.2 The full picture
Having framed the most interesting parameter regions, we are now ready to get the full picture by exploring the more general case of multiple excitations.
@@
To this aim, for computational reasons we shrink our spin chain to $N=8$ and inject $N_{exc}=4$ into the system. We initially inject them all next to each other, as we have seen that this is the condition to size the effect of anisotropy. We also restrict the parameter range to the most significant behavior, as learned from Sec. III.2.1. To sum up, this is for space-homogeneous (e.g. $\nu=100$) collisions happening at smaller rates, high disorder range bandwidth (e.g. $h=10$) and anisotropy as large as $\Delta=2.5$.
@@
We display in Fig. 6(a) the time evolution of the Inverse Ergodicity Ratio (IER) up to enough long time $tJ=1000$, to let the system evolve and potentially reach a delocalized state. We see that for low collision rates ($r_{c}\ll 1$) combined with time-homogeneous collisions, ($\nu\gg 1$), the system struggles to delocalize. In such conditions, the interplay of disorder, anisotropy, and noise stabilizes localized states delaying the onset of delocalization as it is zoomed in the inset. In this case, the height of the plateaus is shorter compared to the case of one and two excitations and complete delocalization is more difficult to reach. Thus, only at very late times does the IER approach its asymptotic value, indicating that the system has completely delocalized.
@@
To gain insight on the underlying spatial behavior, we look in Fig. 6(b) ata color plot for the spread of magnetization over time. We see that for low collision rates and high time-homogeneous collisions, the magnetization piles up for extended periods of time in specific regions of the spin chain, highlighting the system’s localized nature. Anisotropy and disorder favor the pinning of excitations to certain sites. As time progresses, however, collisions eventually disrupt these pinned configurations, leading to gradual spread of magnetization across the chain. Delocalization is however significantly delayed under these conditions, requiring long timescales for the system to be fully delocalized.
@@
Figure 6: Case of multiple excitations. Inverse Ergodicity Ratio (IER) (a) and color plot of the magnetization (b) for a spin chain of $N=8$ sites and $N_{exc}=4$ excitations starting on the left side of the chain, producing a domain-wall initial state, time-homogeneous collisions $\nu=100$, with high disorder range width $h=10$ and anisotropy strength $\Delta=2.5$. (a) IER vs time $tJ$ for different collision rates $r_{c}$ in the interesting regime as in the legend. Inset: Zoom of Fig.(a). (b) Spread of magnetization over time $tJ$ and sites $i$ in the interesting regime of low collision rates, here e.g. with $r_{c}=0.1$. We can overall see that low collision rates ($r_{c}\ll 1$) combined with time-homogeneous collisions ($\nu=100$) lead the system to remain localized for a significant time, prolonged also by the presence of anisotropy and high disorder. Only at very long times the IER approaches its asymptotic value, signaling full delocalization. The simulations were performed with $M=250$ trajectories and timestep $dt=0.02$. Final time is $tJ=1000$.
Before complementing the dynamics characterization with the information that can be extracted from the entanglement entropy, we conclude this section by summarizing in Table I the behaviors described so far.
Parameter Delocalization Localization
Collision rate $r_{c}>1$ $r_{c}\leq 1$
Shape parameter small $\nu$ $\nu>>1$
Disorder strength $h=0.5,1$ $h=5,10$
Table 1: The full picture. Summary of the localization/delocalization regimes varying the collision rate, the shape parameter and the disorder strength.
The table summary is completed by reminding how time non-homogeneous collisions (low $\nu$) and varying anisotropy $\Delta$ influence the dynamics. If we look at regime (3) in Fig. 5, we notice that the data with small collision rate, high disorder strength and $\nu=100$ have the highest complete delocalization time $\tau$. Thus, in general very high value of $\nu$ favors the presence of localized regions. However, when the collision rate increases the shape parameter plays a sub-leading role in localizing/delocalizing the excitations. Also, we notice that the anisotropy $\Delta$ plays a more important role for multiple excitations, enhancing localization when the excitations are next to each other (see also Appendix B).